AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
ISOLDE
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 19:29:09 2025
Author
*
:
Subject
*
:
Header: 3.99 MeV/u and CH2_30+25. Continue after reheat the primary target (GPS) between 02:20 and 02:30 Ender: ==================================================== Run number : 86 Date : 11/11/17 Start time : 02:37:15 Stop time : 04:48:36 Trigger rate (MIDAS) : 2 Blocks written : 87 覧覧覧覧覧覧覧覧覧覧覧覧覧覧覧 Run Scalars (approx rates) 0- Ungated : 3 1- Gated : 3 2- dE : 136k 3- E : 103 4- HI : 432 5- Pulser : 2 6- Clock : 968 覧覧覧覧覧覧覧覧覧覧覧覧覧覧覧 Beam energy (MeV/u) : 3.99 MeV/u Beam intensity (epA) : ~1.5 pA (spread is decreased: 0.7 - 2.2pA) (pictures from 3:44 and 4:35) Target : CH2_30+25 (Pos#6; 490-520 ug/cm2) Ladder position (mm) : 67.26 HI Degrader foil (um) : 10um Al 覧覧覧覧覧覧覧覧覧覧覧覧覧覧覧 del_E voltage (V) : -40 del_E leakage (uA) : 0.63 E voltage (V) : -350 E leakage (uA) : 2.12 E_hi voltage (V) : -14 E_hi leakage (uA) : 2.40 覧覧覧覧覧覧覧覧覧覧覧覧覧覧覧 PA power (Up, positive) (A) : 1.55 PA power (Up, negative) (A) : -0.55 PA power (Down, positive) (A) : 0.8 PA power (Down, negative) (A) : -0.28 覧覧覧覧覧覧覧覧覧覧覧覧覧覧覧 Pressure (IMG1) (mbar) : 8.35 E-06 Pressure (old gauge) (mbar) : 2.6 E-05 Pressure (SEC) (mbar) : 4.6 E-06 覧覧覧覧覧覧覧覧覧覧覧覧覧覧覧 Temp (Chiller) deg : -10.0 Temp (chamber) deg : 5.6 覧覧覧覧覧覧覧覧覧覧覧覧覧覧覧 Comments: 02:14 realized target current was done, but not long. Stopped Run 85 02:18 send protons to beam dump and then called PS to stop them 02:20 in Knob open application: current was set to 0, Reset, turned ON and was increased with medium steps of around 10A. 02:30 target current readout at 438 A (set to 433 A) again. Proton back on target 04:40 Laser quality is decreasing. Watched the bars (green and blue) a while and observed huge fluctuation. Green bar (right) is very low in general, and blue bar (left) is jumping between full and something less of one half. I thought that should be more stabil. 04:46 Events in E_a+E_HI-sepctra: 6 event between 67-129MeV (not sure when it last was reset) --> very low event rate. If the beam is running is mainly checked in MIDAS by looking at trigger(0 and 1) and on refresh update of spectra s3008 (ADC versus channel (s2order) to see new events. n+n-side of detectors still seeing much noise, which is also obvious in the scalar rates. ====================================================
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown